• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 17:46:50 on October 19, 2002

    Entry number 88410

    keyword=Beam Time Accounting
    Hall A

    10/19/2002 0800-1600
    Experiment #: e01020
    Shift Leader: E. Voutier
    Shift Workers: Bogdan W., H. Ibrahim

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    0 0 0 0 0 0 0 0 0
    Beam Available
    but Not in Use
    1 1 1 1 1 1 1 1 8
    Beam Not Available
    or Unacceptable
    0 0 0 0 0 0 0 0 0
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    0 0 0 0 0 0 0 0 0
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0 0
    Unplanned
    Experiment Down
    1 1 1 1 1 1 1 18
    Hall not
    ready reasons

    Comments: Cooldown of target completed around 11:30am. Problem with right arm Q1 fixed and readyaround 2:30pm. Ask to go from restricted to controlled around 2:45pm. After got to controlled access had to make access to fix problem with aerogel adc. Found loose connection at adc. Fixed problem and all ok.

    Data Entry by: M Jones